# Characterization of Deletion/Substitution Channel Capacity for Small Deletion and Substitution Probabilities

### v1

@@ Abstract
In this paper, we consider binary input deletion/substitution channels, which model certain channels with synchronization errors encountered in practice. Specifically, we focus on the regimen of small deletion and substitution probabilities, and by extending an approach developed for the deletion-only channel, we obtain an asymptotic characterization of the channel capacity for independent and identically distributed deletion/substitution channels. We first present an upper bound on the capacity for an arbitrary but fixed number of deletions and substitutions, and then we extend the result to the case of random deletions and substitutions. Our final result is as follows: the i.i.d. deletion/substitution channel capacity is approximately $1-H(p_{d})-H(p_{s})$, where $p_{d}$ is the deletion probability, and $p_{s}$ is the substitution probability.
@@ I Introduction
Channels with insertion, deletions, and substitutions are crucial in modeling certain communication systems with synchronization errors. For instance, when there is a clock mismatch in a wireless communication system, such errors are observed. In addition, bit-patterned media recording channels and DNA data storage channels exhibit such errors. In this paper, our focus is on the special case of channels experiencing bit losses (deletions) and bit errors (substitutions). Specifically, we are interested in independent, identically distributed (i.i.d.) deletions and substitutions and the regime of small deletion and substitution probabilities. Our objective is to characterize the channel capacity for this regime.
Dobrushin established the information stability of channels with memoryless synchronization errors in [1], which is extended by others in [2] to synchronization errors governed by a Markov chain. However, obtaining the exact capacity is very demanding, and even for a channel with i.i.d. deletions, it is still unknown. As a result, the literature has mostly focused on obtaining bounds on the capacity. Gallager’s work on sequential decoding for channels with synchronization errors provides the first lower bound on the capacity of a deletion channel [3], which states that for deletion probabilities of less than $1/2$, i.e., $p_{d}<1/2$, the capacity of i.i.d. binary deletion channel (BDC) is not smaller than that of a binary symmetric channel with a cross-over probability of $p_{d}$; that is
@@
$C_{d,s}\geq 1+p_{d}\log p_{d}+P_{e}\log P_{e}+P_{c}\log P_{c},$ (2)
where $P_{e}=(1-p_{d})p_{s}$ is the probability that the bit is incorrectly transmitted, with $p_{s}$ denoting the substitution probability, and $P_{c}=(1-p_{d})(1-p_{s})$ is the probability that it is correctly transmitted. In the regime of asymptotically small synchronization errors, this can be lower bounded as
@@
We will prove that this bound is asymptotically tight by establishing an upper bound on the capacity whose dominant terms match with it in the asymptotic regime of $p_{d},p_{s}\rightarrow 0$. There are also other attempts to obtain lower bounds on the capacity of channels with synchronization errors. The authors in [4] employ tree codes to obtain lower bounds for insertion/deletion channels. In [5], a lower bound on the BDC capacity is obtained using the jigsaw-puzzle decoding approach, which is further improved in [6].
Only a limited number of upper bounds have been presented for the capacity of channels with synchronization errors. In [7], the authors provide two upper bounds on the BDC capacity: a numerically evaluated one for any arbitrary deletion probability and a simple closed-form one for asymptotically large deletion probabilities ($p_{d}\rightarrow 1$). In [8], numerical upper bounds are provided for deletion/substitution channels by exploiting an auxiliary system where the receiver is provided with side information. The authors extend their results to a non-binary deletion channel in [9], and to more general synchronization error channels with insertions, deletions, and substitutions in [10]. In [11], by converting the capacity problem into maximizing a univariate, real-valued, and often concave function, the authors obtain upper bounds for a general class of repetition channels.
Some works have focused on approximating the capacity of systems with asymptotically small synchronization errors. In [12, 13], the BDC capacity is approximated using series expansions up to the second and third leading terms, respectively, by calculating the dominant terms of the capacity expansion for small deletion probabilities. Similar results are obtained for duplication channels [14] and insertion channels [15].
@@
To obtain this bound, they first obtain lower and upper bounds on the success probability of a guesser that, given the received erroneous sequence, outputs the transmitted sequence along with the deletion pattern. To this end, assuming that the decoding is successful with a certain probability, they obtain bounds on the number of deletion patterns in a certain vicinity of the desired one that can be mistaken with it at the receiver. Finally, combining (4) with Gallager’s lower bound, they show that the BDC capacity approaches $1-H(p_{d})$ as $p_{d}\rightarrow 0$. The authors extend the result to $k$-copy BDC in [17].
In this paper, we build on the result in [16] and extend it to channels with substitution errors as well as deletions. we first provide an upper bound on the capacity of deletion/substitution channels. To this end, we first fix the numbers of deletions and substitutions and assume that the receiver knows them as side information (it knows the numbers of deletions anyway). Next, we cancel the effect of this (side-information) assumption by subtracting its effect. Moreover, combining with the result in [3], we obtain a tight approximation of the capacity when the deletion and substitution probabilities are asymptotically small. Finally, we extend the result to the case of random numbers of deletions and substitutions, and obtain a tight approximation for the capacity in the asymptotic regime of small deletion and substitution probabilities.
@@ II System Model
We consider a deletion/substitution channel where each bit is independently deleted with probability $p_{d}$, and among the transmitted bits, each bit is flipped (substituted) with probability $p_{s}$.
We consider codebook $C\subseteq\{0,1\}^{n}$ consisting of messages of $n$ bits and of size $N=|C|$. We may think of a transmission pattern $A_{t}=a^{t}_{1},a^{t}_{2},...,a^{t}_{n-q_{d}}$ as a strictly increasing subsequence of $[n]:=\{1,2,...,n\}$, representing which bits are transmitted (not deleted), where $q_{d}$ is the number of deletions. Similarly, we denote a substitution pattern $A_{s}$ as a strictly increasing subsequence of $[n-q_{d}]$, $A_{s}=a^{s}_{1},a^{s}_{2},...,a^{s}_{q_{s}}$, where $q_{s}$ is the number of substitutions.
The Deletion/substitution channel can be considered as a concatenation of a deletion channel with probability $p_{d}$ and a substitution channel with substitution probability $p_{s}$, as depicted in Fig. 1. Note that deletion/substitution channel can be considered as the concatenation of a substitution channel with a deletion channel as well, but we found the former presentation easier to work with for our analytical derivations.
Figure 1: A deletion/substitution channel where the input passes first a deletion channel and then a substitution one.
Each realization of the deletion/insertion channel can be represented as a pattern pair $A=(A_{t},A_{s})$. For a string $X\in\{0,1\}^{n}$, $X_{A}$ represents the transmission of $X$ through a deletion/substitution channel with transmission and substitution patterns $A_{t}$ and $A_{s}$, respectively, in an obvious way: the bits of $X$ first get transmitted according to the transmission pattern $A_{t}$, then the remaining bits get flipped with the substitution pattern $A_{s}$.
@@ III Fixing the numbers of deletions and substitutions
We aim to obtain an upper bound on the codebook size (whose logarithm gives us an upper bound on the capacity). We first obtain the upper bound on the codebook size for fixed numbers of deletions and substitutions, assuming that the receiver knows them as side information. Note that the receiver is always aware of the number of deletions as it is equal to the length loss of the received sequence but not the number of substitutions.
To obtain the upper bound on the codebook size, we first obtain the upper and lower bounds on the success probability of a guesser that, given a received sequence, guesses the transmitted codeword as well as the transmission (deletion) and substitution patterns. Then, we compare these upper and lower bounds to obtain an upper bound on the codebook size.
@@
Remark 1
[16, Lemma II.4] Let $\rho$ be a joint distribution over $S\times T$, for finite sets $S$ and $T$, such that the marginal distribution over $S$ is uniform. Let $g:T\rightarrow S$ be a function. Then, $Pr_{(a,b)\sim\rho}[g(b)=a]\leq\frac{|T|}{|S|}$.
@@
Using Remark 1, for fixed numbers of deletions and substitutions, we obtain the following upper bound on the success probability of a general guesser:
@@
To obtain a lower bound on the success probability, we first need to obtain an upper bound on the number of neighboring patterns that can be mistaken with the actual one. To this end, we first upper bound the probability of two transmission/substitution pattern pairs leading to the same output for all possible inputs in the following lemma.
@@
Take any two transmission/substitution pattern pairs, $A=(A_{t},A_{s})$ and $B=(B_{t},B_{s})$. For uniformly random $X\in\{0,1\}^{n}$,
$\begin{split}Pr_{X\in U^{\{0,1\}^{n}}}[X_{A}=X_{B}|(A_{t},A_{s}),&(B_{t},B_{s})]\\ &\begin{cases}\leq 2^{-|\Delta_{t}|},&\text{if }\delta_{s}\subseteq\Delta_{t}\\ =0,&\text{o.w.}\end{cases}\end{split}$ (6)
where $\Delta_{t}$ denotes the discrepancy set between the two transmission patterns and $\delta_{s}$ is the symmetric difference of the two substitution patterns; i.e., $\Delta_{t}=\Delta(A_{t},B_{t})$ and $\delta_{s}=\delta(A_{s},B_{s})$.
@@
In [16], it is shown that in the case of just deletions, the probability of channel outputs being the same is upper bounded as $Pr_{X\in U^{\{0,1\}^{n}}}[X_{A_{t}}=X_{B_{t}}]\leq 2^{-|\Delta_{t}|}$. In the case of just substitutions, two outputs can be the same only if the two substitution patterns are exactly the same. However, that is not the case for the deletion/substitution scenario. In that case, as long as the substitution pattern difference set is a subset of the discrepancy set of the transmission patterns, the two outputs can be the same with the same probability as the two transmission patterns that lead to the same output. So, we have
@@
The expression in 6 can be further upper bounded as
$Pr_{X\in U^{\{}0,1\}^{n}}[X_{A}=X_{B}|(A_{t},A_{s}),(B_{t},B_{s})]\leq 2^{-|\Delta_{t}|}.$ (8)
In the next lemma, we upper bound the number of patterns in a certain neighborhood of a given pattern that can be mistaken with it.
@@
As can be seen in Lemma 6, two patterns can only be mistaken with each other if the set of their substitution differences is a subset of their transmission pattern differences; i.e., $\delta_{s}\subseteq\Delta_{t}$. The number of such patterns, $I$, can be expressed as
$\begin{split}I&=|B||\Delta_{t}|\leq t,|\delta_{s}|\leq s,\delta_{s}\subseteq\Delta_{t}|\\ &\leq\sum_{t^{\prime}=0}^{t}\underbrace{|B||\Delta_{t}|=t^{\prime}|}_{I_{t^{\prime}}}\times\underbrace{|B||\Delta_{t}|=t^{\prime},|\delta_{s}|\leq s,\delta_{s}\subseteq\Delta_{t}|}_{I_{s}}.\end{split}$ (10)
@@
$\begin{split}I_{s}&=|B||\Delta_{t}|=t^{\prime},|\delta_{s}|\leq s,\delta_{s}\subseteq\Delta_{t}|\\ &=\sum_{s^{\prime}=0}^{s}|B||\Delta_{t}|=t^{\prime},|\delta_{s}|=s^{\prime},\delta_{s}\subseteq\Delta_{t}|\\ &\leq\sum_{s^{\prime}=0}^{s}{q_{s}\choose q_{s}-s^{\prime}}{t^{\prime}\choose s^{\prime}},\end{split}$ (12)
@@
Definition: For integers $t$ and $s$, we say $X\in\{0,1\}^{n}$ is $(t-s)$-bad if there exist two patterns $A$ and $B$ such at $|\Delta_{t}|\geq t$ or $|\delta_{s}|\geq s$, and $X_{A}=X_{B}$.
In the following lemma, we provide an upper bound on the number of $(t-s)$-bad sequences.
@@
Fixing the number of deletions and substitutions, the number of $(t-s)$-bad sequences in $X\in\{0,1\}^{n}$, $I_{(t-s)-bad}$, can be upper bounded as
$I_{(t-s)-bad}\leq{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}2^{n-t}.$ (15)
@@
We first obtain the probability of a sequence being $(t-s)$-bad, $P_{(t-s)-bad}$. According to Corollary 8, the probability that two patterns $A$ and $B$ lead to the same output is upper bounded by $2^{-|\Delta_{t}|}$. Since this probability is decreasing in $|\Delta_{t}|$, and using union bound trick, we have
$\begin{split}&P_{(t-s)-bad}\\ &=Pr_{X\in U^{\{0,1\}^{n}}}[\exists A,B|(|\Delta_{t}|\geq t\lor|\delta_{s}|\geq s)\land X_{A}=X_{B}]\\ &\leq\bigcup_{A,B}Pr_{X\in U^{\{0,1\}^{n}}}[X_{A}=X_{B}]\\ &\leq{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}2^{-t}\end{split}$ (16)
@@
Fixing the numbers of deletions and substitutions, suppose there exists a decoding algorithm that succeeds with probability at least $\delta\geq 0$. Then, the size of the codebook satisfies
@@
with
$t=\lceil{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{(n-q_{d})e}{q_{s}}+\log\frac{2}{\delta}\rceil.$
@@
We first assume that the receiver knows the number of deletions and substitutions as side information. We consider a guesser that, given the received sequence $Z_{A}$, outputs the transmitted sequence $Z$ and the pattern pair $A=(A_{t},A_{s})$, both chosen uniformly at random, with a non-negligible probability. More specifically, on the input $Z_{A}$, the guesser first decodes it with the proposed decoder $R(\cdot)$ with success probability $\delta$, then outputs $g(Z_{A})=(R(Z_{A}),A^{\prime})$, where $A^{\prime}$ is the lexicographically first sequence to satisfy $(R(Z_{A}))_{A^{\prime}}=Z_{A}$ if one exists, and the fixed pattern pair $A^{\prime}=(A_{t}^{\prime}=1,2,\cdots,n-q_{d},A_{s}^{\prime}=1,2,...,q_{s})$ otherwise.
We now lower bound the success probability of the guesser. For this, we first consider the probability that the decoding succeeds ($R(Z_{A})=Z$) and the codeword $Z\in\mathcal{C}$ is not $(t-s)$-bad as follows
$\begin{split}&Pr_{Z\in U^{\mathcal{C}}}[R(Z_{A})=Z\wedge Z\text{ is not (t-s)-bad}]\\ &\geq Pr_{Z\in U^{\mathcal{C}}}[R(Z_{A})=Z]+Pr_{Z\in U^{\mathcal{C}}}[Z\text{ is not {(t-s)-bad}}]-1\\ &=\delta-Pr_{Z\in U^{\mathcal{C}}}[Z\text{ is {(t-s)-bad}}]\\ &\geq\delta-\frac{2^{n-t}}{N}{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}\end{split}$ (18)
since $Pr_{Z\in U^{\mathcal{C}}}[Z\text{ is {(t-s)}-bad}]\leq\frac{|\text{{(t-s)}-badsequences}\in\{0,1\}^{n}|}{|\mathcal{C}|}$.
@@
If $Z$ is not $(t-s)$-bad, pattern $B$ gets mixed up with the desired pattern $A$ only when $\Delta_{t}\leq t-1$ and $\delta_{s}\leq s-1$. According to Lemma 9, the number of such patterns is upper bounded by
$\begin{split}&st{q_{s}+s-1\choose q_{s}}{t+s-2\choose s-1}{{q_{d}+t-1}\choose{q_{d}}}{{2q_{d}+t}\choose{2q_{d}+1}}\\ &\leq t^{2}{q_{s}+t-1\choose q_{s}}{2t-2\choose t-1}{{q_{d}+t-1}\choose{q_{d}}}{{2q_{d}+t}\choose{2q_{d}+1}}\\ &\leq t^{2}\left(e\frac{q_{s}+t-1}{q_{s}}\right)^{q_{s}}\left(2e\right)^{t-1}\left(e\frac{q_{d}+t-1}{q_{d}}\right)^{q_{d}}\left(e\frac{2q_{d}+t}{2q_{d}+1}\right)^{2q_{d}+1}\\ &\leq t^{2}\left(2e\right)^{t-1}\left(\frac{5t}{q_{s}}\right)^{q_{s}}\left(\frac{5t}{q_{d}}\right)^{3q_{d}+1}\\ &:=\alpha,\end{split}$ (20)
@@
Conditioned on the decoding succeeding and the codeword not being $(t-s)$-bad, each deletion/substitution pattern is equally likely, and hence, the lexicographically first pattern is correct with probability at least $\alpha^{-1}$. Hence, the total success probability of the guesser is at least

### v2

@@ Abstract
We consider binary input deletion/substitution channels, which model certain channels with synchronization errors encountered in practice. Specifically, we focus on the regime of small deletion and substitution probabilities, and by extending an approach developed for the deletion-only channel, we obtain an asymptotic characterization of the channel capacity for independent and identically distributed deletion/substitution channels. We first present an upper bound on the capacity for arbitrary but fixed numbers of deletions and substitutions, and then we extend the result to the case of random deletions and substitutions. Our final result is as follows: The i.i.d. deletion/substitution channel capacity is approximately $1-H(p_{d})-H(p_{s})$, for $p_{d},p_{s}\approx 0$, where $p_{d}$ is the deletion probability, and $p_{s}$ is the substitution probability.
@@ I Introduction
Channels with insertions, deletions, and substitutions are crucial in modeling certain communication systems with synchronization errors. For instance, when there is a clock mismatch in a wireless communication system, such errors are observed. In addition, bit-patterned media recording channels and DNA data storage channels exhibit such errors. In this paper, our focus is on the special case of channels experiencing bit losses (deletions) and bit errors (substitutions). Specifically, we are interested in independent, identically distributed (i.i.d.) deletions and substitutions at the regime of small deletion and substitution probabilities. Our objective is to characterize the channel capacity for this asymptotic regime.
Dobrushin established the information stability of channels with memoryless synchronization errors in [1], which is extended in [2] to synchronization errors governed by a Markov chain. However, obtaining the exact capacity has proved very difficult. Even for a channel with only i.i.d. deletions, the capacity is still unknown. As a result, the literature has focused on obtaining bounds on the capacity. Gallager’s work on sequential decoding for channels with synchronization errors provides the first lower bound on the capacity of a deletion channel [3], which states that for deletion probabilities of less than $1/2$, i.e., $p_{d}<1/2$, the capacity of i.i.d. binary deletion channel (BDC) is not smaller than that of a binary symmetric channel with a cross-over probability of $p_{d}$; that is,
@@
$C_{d,s}\geq 1+p_{d}\log p_{d}+p_{s}\log p_{s}+p_{c}\log p_{c},$ (2)
where $p_{s}$ is the substitution probability, and $p_{c}=1-p_{d}-p_{s}$ is the probability of correct transmission. In the regime of asymptotically small synchronization errors, this can be lower bounded as
@@
We will prove that this bound is asymptotically tight by establishing an upper bound on the capacity whose dominant terms match it in the asymptotic regime of $p_{d},p_{s}\rightarrow 0$. There are also other attempts to obtain lower bounds on the capacity of channels with synchronization errors. The authors in [4] employ tree codes to obtain lower bounds for insertion/deletion channels. In [5], a lower bound on the BDC capacity is obtained using the jigsaw-puzzle decoding approach, which is further improved in [6].
Only a limited number of upper bounds have been presented for the capacity of channels with synchronization errors. In [7], the authors provide two upper bounds on the BDC capacity: a numerically evaluated one for any arbitrary deletion probability and a simple closed-form one. In [8], numerical upper bounds are provided for deletion/substitution channels by exploiting an auxiliary system where the receiver is provided with side information. The authors extend their results to a non-binary deletion channel in [9], and to more general synchronization error channels with insertions, deletions, and substitutions in [10]. The authors of [11] improve the BDC lower bound in [6] as well as the upper bound in [8]. In [12], by converting the capacity problem into maximizing a univariate, real-valued, and often concave function, the authors obtain upper bounds for a general class of repetition channels.
Several works have focused on approximating the capacity of synchronization error channels for the asymptotic regime of small synchronization errors. In [13, 14], the BDC capacity is approximated using series expansions up to the second and third leading terms, respectively, by calculating the dominant terms of the capacity expansion for small deletion probabilities. Similar results are obtained for duplication channels [15] and insertion channels [16].
@@
To obtain this bound, they first obtain lower and upper bounds on the success probability of a guesser that, given the received erroneous sequence, outputs the transmitted sequence along with the deletion pattern. To this end, assuming that the decoding is successful with a certain probability, they obtain bounds on the number of deletion patterns in a certain vicinity of the desired one that can be mistaken for it at the receiver. Finally, combining (4) with Gallager’s lower bound, they show that the BDC capacity approaches $1-H(p_{d})$ as $p_{d}\rightarrow 0$. The authors extend the result to the $k$-copy BDC in [18].
In this paper, we adapt the general methodology in [17] and extend it to channels with substitution errors as well as deletions. We first provide an upper bound on the capacity of deletion/substitution channels. To this end, we first fix the numbers of deletions and substitutions and assume that the receiver knows these numbers as side information (it knows the number of deletions anyway due to the received sequence length), and obtain an upper bound on the capacity, which is also an upper bound on the capacity of the case without side information since side information cannot decrease capacity. Combining with the result in [3], we obtain a tight approximation of the capacity when the deletion and substitution probabilities are asymptotically small. Finally, we extend the result to the case of random numbers of deletions and substitutions, and obtain a tight approximation for the capacity in the asymptotic regime of small deletion and substitution probabilities.
@@ II System Model
We consider a deletion/substitution channel, where each bit is independently deleted with probability $p_{d}$, flipped (substituted) with probability $p_{s}$, or correctly transmitted with probability $1-p_{d}-p_{s}$. We consider a codebook $C\subseteq\{0,1\}^{n}$ consisting of messages of $n$ bits and of size $N=|C|$. We may think of a transmission pattern $A_{t}=a^{t}_{1},a^{t}_{2},...,a^{t}_{n-q_{d}}$ as a strictly increasing subsequence of $[n]:=\{1,2,...,n\}$, representing which bits are transmitted (not deleted), where $q_{d}$ is the number of deletions. Similarly, we denote a substitution pattern $A_{s}$ as a strictly increasing subsequence of $[n-q_{d}]$, $A_{s}=a^{s}_{1},a^{s}_{2},...,a^{s}_{q_{s}}$, where $q_{s}$ is the number of substitutions.
A deletion/substitution channel with deletion probability $p_{d}$ and substitution probability $p_{s}$ can be decomposed as a concatenation of a deletion channel with deletion probability $p_{d}$ followed by a substitution channel with substitution probability $p_{s}^{\prime}=\frac{p_{s}}{1-p_{d}}$, as depicted in Fig. 1, which we found to be easier to work with for our analytical derivations.
Figure 1: Decomposition of a deletion/substitution channel with deletion probability $p_{d}$ and substitution probability $p_{s}$ as a concatenation of a deletion channel with deletion probability $p_{d}$ followed by a substitution channel with substitution probability $p_{s}^{\prime}=\frac{p_{s}}{1-p_{d}}$.
Each realization of the deletion/insertion channel can be represented as a pattern pair $A=(A_{t},A_{s})$. For a string $X\in\{0,1\}^{n}$, $X_{A}$ represents the transmission of $X$ through a deletion/substitution channel with transmission and substitution patterns $A_{t}$ and $A_{s}$, respectively, in the obvious way: the bits of $X$ are deleted according to the transmission pattern $A_{t}$, then the remaining bits are flipped with the substitution pattern $A_{s}$.
@@ III Fixing the numbers of deletions and substitutions
We aim to obtain an upper bound on the size of a uniquely decodable codebook, whose logarithm gives us an upper bound on the capacity. We first obtain the upper bound on the codebook size for fixed numbers of deletions and substitutions, assuming that the receiver knows them as side information. Note that the receiver is always aware of the number of deletions, as it is equal to the length loss of the received sequence, but not the number of substitutions.
To obtain the upper bound on the codebook size, we first obtain upper and lower bounds on the success probability of a guesser that, given a received sequence, guesses the transmitted codeword as well as the transmission (deletion) and substitution patterns. Then, we compare these upper and lower bounds to obtain an upper bound on the codebook size.
@@
Lemma 1 ([17])
Let $\rho$ be a joint distribution over $S\times T$, for finite sets $S$ and $T$, such that the marginal distribution over $S$ is uniform. Let $g:T\rightarrow S$ be a function. Then, for $(a,b)$ following the joint distribution $\rho$, $Pr[g(b)=a]\leq\frac{|T|}{|S|}$.
@@
Using Lemma 1, for fixed numbers of deletions and substitutions, we obtain the following upper bound on the success probability of a general guesser:
@@
To obtain a lower bound on the success probability, we first need to obtain an upper bound on the number of neighboring patterns that can be mistaken for the actual one. To this end, we first upper bound the probability of two transmission/substitution pattern pairs leading to the same output for all possible inputs in the following lemma.
@@
Take any two transmission/substitution pattern pairs, $A=(A_{t},A_{s})$ and $B=(B_{t},B_{s})$. For uniformly random $X\in\{0,1\}^{n}$, $Pr_{X\in U^{\{0,1\}^{n}}}[X_{A}=X_{B}|(A_{t},A_{s}),(B_{t},B_{s})]\leq 2^{-|\Delta_{t}|}$ if $\delta_{s}\subseteq\Delta_{t}$, and zero otherwise, where $\Delta_{t}$ denotes the discrepancy set between the two transmission patterns and $\delta_{s}$ is the symmetric difference of the two substitution patterns [^1] Note that $\Delta_{t}$ is defined over the indices of the elements of the two sequences, not on their values.; i.e., $\Delta_{t}=\Delta(A_{t},B_{t})$ and $\delta_{s}=\delta(A_{s},B_{s})$.
@@
In [17], it is shown that in the case of only deletions, the probability of channel outputs being the same is upper bounded as $Pr_{X\in U^{\{0,1\}^{n}}}[X_{A_{t}}=X_{B_{t}}]\leq 2^{-|\Delta_{t}|}$. In the case of only substitutions, two outputs can be the same only if the two substitution patterns are exactly the same. However, that is not the case for the deletion/substitution scenario. In that case, as long as the substitution pattern difference set is a subset of the discrepancy set of the transmission patterns, the two outputs can be the same with the same probability as the two transmission patterns that lead to the same output. So, we have
@@
The expression in Lemma 1 can be further upper bounded as $Pr_{X\in U^{\{}0,1\}^{n}}[X_{A}=X_{B}|(A_{t},A_{s}),(B_{t},B_{s})]\leq 2^{-|\Delta_{t}|}$.
In the next lemma, we upper bound the number of patterns in a certain neighborhood of a given pattern that can be mistaken for it.
@@
From Lemma 1, two patterns can only be mistaken with each other if the set of their substitution differences is a subset of their transmission pattern differences; i.e., $\delta_{s}\subseteq\Delta_{t}$. The number of such patterns, $I$, can be expressed as
$\begin{split}&I=|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|\leq t,|\delta_{s}|\leq s\}|\\ &\leq\sum_{t^{\prime}=0}^{t}\underbrace{|\{B||\Delta_{t}|=t^{\prime}\}|}_{I_{t^{\prime}}}\times\underbrace{|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|\leq s\}|}_{I_{s}}.\end{split}$ (8)
@@
$\begin{split}I_{s}&=|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|\leq s\}|\\ &=\sum_{s^{\prime}=0}^{s}|\{B|\delta_{s}\subseteq\Delta_{t},|\Delta_{t}|=t^{\prime},|\delta_{s}|=s^{\prime}\}|\\ &\leq\sum_{s^{\prime}=0}^{s}{q_{s}\choose q_{s}-s^{\prime}}{t^{\prime}\choose s^{\prime}},\end{split}$ (10)
@@
Definition: For integers $t$ and $s$, we say that $X\in\{0,1\}^{n}$ is $(t,s)$-bad if there exist two patterns $A$ and $B$ such that $|\Delta_{t}|\geq t$ or $|\delta_{s}|\geq s$, and $X_{A}=X_{B}$.
In the following lemma, we provide an upper bound on the number of $(t,s)$-bad sequences.
@@
Fixing the numbers of deletions and substitutions, the number of $(t,s)$-bad sequences in $X\in\{0,1\}^{n}$, $I_{(t,s)-bad}$, can be upper bounded as
$I_{(t,s)-bad}\leq{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}2^{n-t}.$ (13)
@@
We first obtain the probability that a sequence is $(t,s)$-bad, $P_{(t,s)-bad}$. According to Corollary 2, the probability that two patterns $A$ and $B$ lead to the same output is upper bounded by $2^{-|\Delta_{t}|}$. Since this probability is decreasing in $|\Delta_{t}|$, and using union bound trick, we have
$P_{(t,s)-bad}$ $=\!Pr_{X\in U^{\{0,1\}^{n}}}[\exists A,B|(|\Delta_{t}|\!\geq t\lor\!|\delta_{s}|\geq s)\!\land\!X_{A}\!\!=\!\!X_{B}]$
$\leq\sum_{A,B}Pr_{X\in U^{\{0,1\}^{n}}}[X_{A}=X_{B}]$
($\leq{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}2^{-t}$ (14)
@@
Fixing the numbers of deletions and substitutions, suppose that there exists a decoding algorithm that succeeds with probability at least $\delta\geq 0$. Then, the size of the codebook satisfies
@@
with $t=\lceil{3q_{d}}\log\frac{ne}{q_{d}}+{3q_{s}}\log\frac{(n-q_{d})e}{q_{s}}+\log\frac{2}{\delta}\rceil.$
@@
We first assume that the receiver knows the numbers of deletions and substitutions as side information. We consider a guesser that, given the received sequence $Z_{A}$, outputs the transmitted sequence $Z$ and the pattern pair $A=(A_{t},A_{s})$, both chosen uniformly at random, with a non-negligible probability. More specifically, on the input $Z_{A}$, the guesser first decodes it with the proposed decoder $R(\cdot)$ with success probability $\delta$, then outputs $g(Z_{A})=(R(Z_{A}),A^{\prime})$, where $A^{\prime}$ is the lexicographically first sequence to satisfy $(R(Z_{A}))_{A^{\prime}}=Z_{A}$ if one exists, and the fixed pattern pair $A^{\prime}=(A_{t}^{\prime}=1,2,\cdots,n-q_{d},A_{s}^{\prime}=1,2,...,q_{s})$ otherwise.
We now lower bound the success probability of the guesser. For this, we first consider the probability that the decoding succeeds ($R(Z_{A})=Z$) and the codeword $Z\in\mathcal{C}$ is not $(t,s)$-bad as follows
[…]
$\geq\delta-\frac{2^{n-t}}{N}{n\choose q_{d}}^{2}{n-q_{d}\choose q_{s}}^{2}$ (16)
since $Pr_{Z\in U^{\mathcal{C}}}[Z\text{ is {(t,s)}-bad}]\leq\frac{|\text{{(t,s)}-badsequences}\in\{0,1\}^{n}|}{|\mathcal{C}|}$.
@@
If $Z$ is not $(t,s)$-bad, pattern $B$ gets mixed up with the desired pattern $A$ only when $\Delta_{t}\leq t-1$ and $\delta_{s}\leq s-1$. According to Lemma 7, the number of such patterns is upper bounded by
$st{q_{s}+s-1\choose q_{s}}{t+s-2\choose s-1}{{q_{d}+t-1}\choose{q_{d}}}{{2q_{d}+t}\choose{2q_{d}+1}}$
$\leq t^{2}{q_{s}+t-1\choose q_{s}}{2t-2\choose t-1}{{q_{d}+t-1}\choose{q_{d}}}{{2q_{d}+t}\choose{2q_{d}+1}}$
$\leq t^{2}\left(e\frac{q_{s}+t-1}{q_{s}}\right)^{q_{s}}\left(2e\right)^{t-1}\left(e\frac{q_{d}+t-1}{q_{d}}\right)^{q_{d}}\left(e\frac{2q_{d}+t}{2q_{d}+1}\right)^{2q_{d}+1}$
$\leq t^{2}\left(2e\right)^{t-1}\left(\frac{5t}{q_{s}}\right)^{q_{s}}\left(\frac{5t}{q_{d}}\right)^{3q_{d}+1}$
$:=\alpha,$ (18)
@@
Conditioned on the decoding succeeding and the codeword not being $(t,s)$-bad, each deletion/substitution pattern is equally likely, and hence, the lexicographically first pattern is correct with probability at least $\alpha^{-1}$. Hence, the total success probability of the guesser is at least